Two ladybugs sit on a rotating disk, as shown in t
Answer #1

a) angular velocity is same for both ladybugs

b) linear speed ratio = v2 / v1 = w r2 / w r1 = r2 / r1 = 2 r / r = 2

c) radial acceleration ratio = w^2 r2 / w^2 r1 = r2 / r1 = 2

d) direction of angular velocity of bug2 = + z

e) direction of tangential acceleration = - y